OpenStudy (anonymous):

Solve the differential equation y' = 0.3y given the condition that y(0) = 9.

zepdrix (zepdrix):

So uhh looks like this one is separable if you just wanna do it that way,\[\Large\rm \frac{dy}{dx}=0.3y\qquad\to\qquad \frac{dy}{y}=0.3dx\]Understand how to integrate each side?

OpenStudy (anonymous):

Okay so what am I supposed to integrate?

zepdrix (zepdrix):

There is a differential on each side, so integrate each side with respect to it's given variable. \[\Large\rm \int\limits \frac{dy}{y}=\int\limits 0.3dx\]

OpenStudy (anonymous):

I got 3x/10 +C

OpenStudy (anonymous):

I don't think it is right though because that is not one of the answers I have.

zepdrix (zepdrix):

I don't understand what you're saying... You got that from integrating the right side? Or the left..? Or what..?

OpenStudy (anonymous):

the right side

zepdrix (zepdrix):

Oh .3x+C, I see. And what about integrating the left side? remember your integral for 1/x?

OpenStudy (anonymous):

the integral for 1/x is ln(x)

OpenStudy (anonymous):

right?

zepdrix (zepdrix):

yes, good. So integrating each side you should get:\[\Large\rm \ln y+c_1=0.3x+c_2\]The constants are arbitrary so let's combine them, subtract c1 from each side, absorb it into the other c,\[\Large\rm \ln y=0.3x+C\]

zepdrix (zepdrix):

Before solving for y, it might be a good idea to use your initial data to find your C value.

zepdrix (zepdrix):

\[\Large\rm y(0)=9\quad means\quad y=9~when~x=0\] Plug it in and do some stuff. \[\Large\rm \ln 9=0.3(0)+C\]

OpenStudy (anonymous):

I am finding C right now. Right?

OpenStudy (anonymous):

k got 2.197

zepdrix (zepdrix):

Uhhh yes, but let's leave it as ln9.

zepdrix (zepdrix):

So your goal is to get this equation, \[\Large\rm \ln y=0.3x+\ln9\]in terms of y=stuff. See how the log is causing a problem?

OpenStudy (anonymous):

yes how do we fix tht problem?

zepdrix (zepdrix):

Exponentiate each side, Rewrite each side as an exponent with a base of e,\[\Large\rm e^{\ln y}=e^{0.3x+\ln9}\]

zepdrix (zepdrix):

Exponential and log are inverse functions of one another, they simply "undo" one another,\[\Large\rm y=e^{0.3x+\ln9}\]But it can be simplified a little further,

zepdrix (zepdrix):

Applying rules of exponents to the right side,\[\Large\rm y=e^{\ln9}\cdot e^{0.3x}\]Then you can apply the exponent/log tricky again.

OpenStudy (anonymous):

=9e^0.3x

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

Awesome! Thanks!